Henry just installed a new rectangular pool in his backyard. The length of the pool is 24 feet more than the width. The perimeter is 88 feet. Find the length and width.
Answer by farbonara(13)

Rectangle is made up of 2 lengths and 2 widths
Excess length=24x2=48
Finding 4 equal parts=88-48=40
Finding 1 equal part=40÷4=10
Width=10
Length=10+24=34